OpenStudy (anonymous):

Three times the first number plus the second number twice the third number is five. If three times the second number is subtracted from the sum of the first number and three times the third number, the result is two. If the third number is subtracted from two times the first number and three times the second number, the result is one. Find the numbers.

OpenStudy (anonymous):

The three equations I came up with from the word problem was: 3x+y+2z=5 x+3y-3y=2 2x+3y-z=1 I believe I am right up to this point

OpenStudy (bahrom7893):

ok solve them..

OpenStudy (anonymous):

So subtract twice the second equation from the first one? and then subtract twice the second ,from the third?

OpenStudy (anonymous):

subtract 3 times the second from the first sorry

OpenStudy (bahrom7893):

yea

OpenStudy (bahrom7893):

what do u get?

OpenStudy (anonymous):

messed up on second equation

OpenStudy (bahrom7893):

wait.. let me check..

OpenStudy (anonymous):

should be x+3z-3y

OpenStudy (anonymous):

so 10y-7z=-1 for first work out

OpenStudy (bahrom7893):

actually there's an easier way i think..

OpenStudy (anonymous):

and 9y-7z=-3

OpenStudy (bahrom7893):

x-3y+3z=2 2x+3y- z=1 -x+4z=1

OpenStudy (anonymous):

right I did switch it to that

OpenStudy (bahrom7893):

ok good.. now

OpenStudy (anonymous):

hmmm see if I come up with the same

OpenStudy (anonymous):

I get y=2

OpenStudy (anonymous):

z=3

OpenStudy (anonymous):

x=-1

OpenStudy (anonymous):

see if that works lol and thank you for the help

OpenStudy (bahrom7893):

that's all good.

OpenStudy (bahrom7893):

haha i barely did anything, but here's to check.. this is right

OpenStudy (anonymous):

hmmm not coming out right

OpenStudy (anonymous):

well maybe it is right lol

OpenStudy (bahrom7893):

8-3=5 first one is satisfied

OpenStudy (anonymous):

that matrices?

OpenStudy (bahrom7893):

yea it's a shortcut i used haha..

OpenStudy (anonymous):

haven't learned that yet but looks easier lol

OpenStudy (bahrom7893):

9-7 = 2 works too..

OpenStudy (bahrom7893):

6-5=1 yup it's all good just double checked on paper

OpenStudy (anonymous):

awesome... thank you

OpenStudy (bahrom7893):

np